OpenStudy (anonymous):

10. A bag contains 8 red marbles and 12 blue marbles. One marble is chosen at random and returned to the bag. Then, a second marble is chosen at random. What is the probability that the first marble chosen is blue and the second marble chosen is red?

OpenStudy (anonymous):

\[\frac{12}{20}\times \frac{8}{20}\]

OpenStudy (anonymous):

\[ \frac{12}{20}\times \frac{8}{20} =\frac{6}{25}\]